# Electronic structure and transport in materials with flat bands: 2D materials and quasicrystals

### v1

@@ 2.3.3 Low-energy flat bands in magic angle twisted bilayer graphene
In the domain of small angles ($\theta<2^{\circ}$), not all TBGs have low-energy flat bands. For certain rotation angles, the so-called “magic angles[93], the low-energy states, i.e., the states with energy close to the Fermi energy for the neutral system, are strongly localized in the AA stacking regions of the moiré pattern (Fig. 6(a)) and form flat bands [90, 92, 93]. [^2] In this section we focus on electronic properties of perfect TBGs, i.e., without defect, at the magic angle. The effect of resonant defects on TBG’s conductivity with various rotation angles $\theta$ is presented in more detail in Sec. 3.2.3. A specificity of this localization by the TBG moiré is that it contains very few electronic states: only four bands per moiré (Fig. 7), i.e., four states over more than 10000 $p_{z}$ orbitals. Moreover, these states are located on large areas of the moiré (AA stacking regions), i.e., on thousands of orbitals. The spatial extent of the corresponding wave packet is thus much larger than the interatomic distance. By means of a simple quantization argument [94] in the AA zone of the moiré, it is possible to find a simple condition for the confinement of Dirac electrons by the moiré pattern that allows us to predict the magic angle series found numerically from a continuum model [93]:
@@
In 2018, it was experimentally proven that this electronic localization by the moiré geometry can induce strong electronic correlations and a superconducting state for some fillings of the four flat bands [102, 103, 245]. This new and exciting condensed matter physics joins in many aspects recent studies under the bicrystallography [101, 100], i.e., systems for which the atomic structure results from commensurable or non-commensurable modulation between two periodicities (or pseudo-periodicity), Eq. (18), of very different length (interatomic-scale periodicity and moiré-scale periodicity). In those complex systems, the consequences of crystallographic symmetries on the electronic properties are not yet well understood.
@@
Before detailing the electronic properties of flat bands in TBGs, it is interesting to note that these flat bands are not systematic when there is a moiré pattern. For example, we have not been able to obtain similar flat bands by calculations using the TB model for bilayer graphene with the similar moiré size but resulting from biaxial strain of one of the two layers (hetero-biaxial strain) without atomic relaxation. On the other hand, a very small hetero-biaxial strain (with atomic relaxation) can produce a moiré of nearly four million atoms [243]. It leads to strong atomic relaxation with the AA regions almost disappearing, the appearance of swirls that delimit large AB stacking zones, and new electronic confinements in and between these zones.
@@
Obviously, as shown by the calculation of the intraband velocity $V_{TBG}({\rm K})$ at K (Fig. 6(d)), the magic angle value depends on the TB model: $\theta_{1^{st}magic}\approx 1.2^{\circ}$ and $\approx 0.9^{\circ}$ for $V_{MLG}=0.79\cdot 10^{6}$ ms-1 and $1.09\cdot 10^{6}$ ms-1 for the SK-TB model, respectively. It should also be noted that $V_{\rm K}$ is very low over a fairly wide range of angles of the order of $\Delta\theta\approx 0.2^{\circ}$. This is well confirmed by the fact that the bands are fairly flat, without ever being strictly flat, for a similar range of angles around the so-called magic angle (Fig. 7). In this range of angles, other effects such as uniaxial heterostrain have a very strong effect on the bands [104, 105, 106] (see Sec. 2.3.5).
@@
Figure 7 shows typical examples of the four flat bands for different angles around the first magic angle. Depending on the value of the angle, these bands can be isolated by minigaps or not strictly isolated, although in the latter case they are almost isolated by pseudogaps with a DOS very close to zero. So, from figure 7, for $\theta=1.248^{\circ}$, $1.203^{\circ}$ and $1.085^{\circ}$ the four flat bands are not separated by minigaps in the rest of the spectrum, whereas this is the case for $\theta=1.297^{\circ}$. It is sometimes argued in the literature that atomic relaxation is necessary to obtain these minigaps, but our results show that this is not the case, and that the shape of the bands vary strongly with angle, while remaining fairly flat overall. These bands are never perfectly flat, and it is not easy to define a “good” criterion for determining the most flat bands. Usually, as previously discussed, the intraband velocity of the Dirac cone is used, and thus the magic angle is obtained when this velocity goes to zero. However, other criteria such as bandwidth may also be relevant in some cases and may give a different $\theta$ value for the magic angles.
@@
The local DOS in the AA zone confirms the main localization of the flat-band states in the AA zone. Detailed analysis of these flat-band states, in particular by the participation ratio calculation [29], shows that they are all equally distributed on the two layers, whereas for intermediate angles, the distribution of each state is mainly in one of the two layers (see participation ratio in Fig. 6(g)).
@@
Figure 8: Electronic properties of asymmetrically doped TBGs. TB calculation with an on-site energy difference $\Delta\epsilon$ between the two layers. (a) Sketch of the Dirac cones and total DOS of a doped rotated bilayer. (b) Energy difference $\Delta E_{D}$ between the energies of two Dirac cones in TBG versus $\Delta\epsilon$. For $\Delta\epsilon$ small, $\Delta E_{D}\approx\Delta\epsilon$ for large angles whereas $\Delta E_{D}<\Delta\epsilon$ for small angles. (c) Band velocity at the K point versus $\Delta\epsilon$. The vertical dashed lines show the value $\Delta\epsilon$ for which $\Delta E_{D}(\Delta\epsilon)$ is maximum (panel (b)). (d), (e) Bands in (25,26) TBG for $\Delta\epsilon=0$ and $\Delta\epsilon=0.2$ eV, respectively. Dashed lines are the monolayer bands. (f) Zoom of panel (e). (g) Average band velocity $V_{TBG}$ versus energy $E$ for (25,26) TBG for different $\Delta\epsilon$ values ($V_{TBG}$ is the Boltzmann velocity computed from the velocity operator and Eq. (72)). TBG: $(5,9)$ $\theta=18.73^{\circ}$, $(2,3)$ $\theta=13.17^{\circ}$, $(3,4)$ $\theta=9.43^{\circ}$, $(6,7)$ $\theta=5.08^{\circ}$, $(8,9)$ $\theta=3.89^{\circ}$, $(12,13)$ $\theta=2.13^{\circ}$, $(25,26)$ $\theta=1.30^{\circ}$. SK-TB model is the one which results in an intraband monolayer velocity $V_{MLG}=0.79\cdot 10^{6}$ ms-1 (Sec. 2.3.2). TBGs are rigid, i.e., without atomic relaxation. From [29].
[…]
@@ 2.3.4 Quantum diffusion
Figure 9: Diffusivity $\mathcal{D}$ versus scattering time $\tau$ in graphene, TBGs (5,9), (6,7), and (25,26), with rotation angle $\theta=18.73^{\circ}$, $5.08^{\circ}$ and $1.297^{\circ}$, respectively. For (25,26) TBG the Boltzmann (B) term and non-Boltzmann (NB) term are shown, $\mathcal{D}(E_{F},\tau)=\mathcal{D}_{\rm B}(E_{F},\tau)+\mathcal{D}_{\rm NB}(E_{F},\tau)$ (see Eq. (75)). These calculations are performed with an energy broadening of 10 meV around $E_{F}$. For (25,26) TBG that corresponds mainly to the flat bands (Fig. 7(a)), i.e., the minimum of $V_{\mathrm{B}}$ (Fig. 8(g)). (a) $\Delta\epsilon=0$, $E_{F}=0.12$ eV (see bands and DOS in Fig.7(a)), and (b) $\Delta\epsilon=0.2$ eV, $E_{F}=-0.1$ eV. $\Delta\epsilon$ the energy difference between the on-site energy in the two layers (Sec. 2.3.3). SK-TB model is the one which results in an intraband monolayer velocity $V_{MLG}=0.79\cdot 10^{6}$ ms-1 (Sec. 2.3.2). TBGs are rigid, i.e., without atomic relaxation. Numerical method see A.2.1. From [29].
@@
In this section, we describe the consequences of the particular localization of TBG flat band states on their quantum diffusion properties [29, 124, 128, 129, 132]. Here we are not considering the defect effect, which produces particular states, and will be studied in the next section (Sec. 3.2.3). We are only considering quantum diffusion in the defect-free medium, and we are treating the effect of disorder by a scattering time $\tau$. As explained in section 2.1 (see also A.2), the diffusivity $\mathcal{D}$ at every energy $E$ and scattering time $\tau$ is the sum of a Boltzmann term, $\mathcal{D}_{\rm B}(E_{F},\tau)=V_{\mathrm{B}}^{2}\tau$, and a non-Boltzmann term $\mathcal{D}_{\rm NB}$. For large $\tau$, $\mathcal{D}_{\rm NB}$ decreases when $\tau$ increases, and $\mathcal{D}_{\rm NB}\rightarrow 0$ when $\tau\rightarrow+\infty$. Thus in crystals, $\mathcal{D}\approx\mathcal{D}_{\rm B}$ when $\tau\rightarrow+\infty$.
@@
The diffusivity, calculated by diagonalization (A.2), for monalayer graphene and several cases of TBGs are presented in Fig. 9 for different $E_{F}$ values and for doped or undoped bilayers. For graphene and bilayers with large and intermediate rotation angles, $\mathcal{D}\approx\mathcal{D}_{\rm B}$ at every energy. The only effect of the non-Boltzmann term is a change in the slope of $\mathcal{D}(\tau)$ at scattering time $\tau\approx\hbar/E$ as explained in the appendix of Ref. [29]. Eventually at small scattering time, $\tau\ll\hbar/E$, the interband transition between the two bands of each Dirac cone contributes significantly. In the case of graphene, with a first-neighbor coupling Hamiltonian, the non-Boltzmann term $\mathcal{D}_{\rm NB}$ is equal to the Boltzmann term $\mathcal{D}_{\rm B}$ and then $\mathcal{D}=2\mathcal{D}_{\rm B}$ for $\tau\ll\hbar/E$ [29]. This effect is related to the phenomenon of jittery motion, also called Zitterbewegung [239], which is important in the optical conductivity. In graphene and bilayers with large rotation angles, it occurs for very small scattering time values that are too small to be significant experimentally. However, in rotated bilayers with very small rotation angle $\theta$ close to the first magic angle, for states at energy where velocity is very small (i.e., energies close to the Dirac energy), the Boltzmann terms in the transport coefficients (see equations (8) and in A.2) decrease and the non-Boltzmann terms become significant. As we have seen in the case of quasiperiodic or approximant structures, this has a strong effect on the charge carrier diffusion regime and therefore on the effect of disorder.
@@
For instance, Fig. 9 shows that for (25,26) bilayers ($\theta=1.297^{\rm o}$), the diffusivity $\mathcal{D}$ is strongly affected by the non-Boltzmann term for realistic $\tau$ values [246]. It results in a smaller diffusivity with respect to the case of graphene, that is not at all ballistic; indeed $\mathcal{D}$ is not at all linear as a function of $\tau$ for realistic values of the scattering time, $\tau\approx 10^{-14}$$10^{-13}$s. Even more, at charge neutrality (i.e., $E_{F}=0.012$ eV for (25,26) TBG see Fig. 7(a)) and without asymmetric doping between the two layers (Fig. 9(a)), diffusivity decreases from $\mathcal{D}=6\cdot 10^{-3}$ m2s-1 to $3\cdot 10^{-3}$ m2s-1, as $\tau$ increases from $\tau\approx 4\cdot 10^{-15}$ s and $10^{-13}$ s. This decrease in conductivity by a factor of 2 due to increased disorder, which may be static (defects) or dynamic (temperature) clearly shows that disorder can lead to an increase in conductivity of the flat-band states, as expected in the sub-diffusive regime (see Sec. 2.2.2). Since flat-band states are localized over large areas of space (several hundreds atoms in AA zones) and their band velocity is very low, it is clear that these states satisfy the SVR condition (Eq. (6)). It also shows indirectly that flat-band states are fragile and can be destroyed by disorder. Recent real space calculations of conductivity using Kubo’s formula have shown such behavior for a magic-angle TBG in the presence of Anderson disorder [129]. In asymmetric doped TBG (Fig. 9(b)), similar effects occur at energies with small Boltzmann velocity (Fig. 8(g)). This regime, called small velocity regime (SVR), where non-Boltzmann terms dominate transport properties, has already been observed in systems with complex atomic structure such as quasicrystals and complex metallic alloys (Sec. 2.2.1). Let us recall that SVR is reached when the mean free path $V\tau$ of charge carriers is smaller than the spatial extent $L_{w}$ of the corresponding wave packet for large time $t$, i.e., $t\gtrsim\tau$.

### v2

@@ 2.3.3 Low-energy flat bands in magic angle twisted bilayer graphene
In the domain of small angles ($\theta<2^{\circ}$), not all TBGs have low-energy flat bands. For certain rotation angles, the so-called “magic anglesBistritzer11 , the low-energy states, i.e., the states with energy close to the Fermi energy for the neutral system, are strongly localized in the AA stacking regions of the moiré pattern (Fig. 6(a)) and form flat bands Trambly10 , SuarezMorell10 , Bistritzer11 . [^2] In this section we focus on electronic properties of perfect TBGs, i.e., without defect, at the magic angle. The effect of resonant defects on TBG’s conductivity with various rotation angles $\theta$ is presented in more detail in Sec. 3.2.3. A specificity of this localization by the TBG moiré is that it contains very few electronic states: only four bands per moiré (Fig. 7), i.e., four states over more than 10000 $p_{z}$ orbitals. Moreover, these states are located on large areas of the moiré (AA stacking regions), i.e., on thousands of orbitals. The spatial extent of the corresponding wave packet is thus much larger than the interatomic distance. By means of a simple quantization argument Trambly12 in the AA zone of the moiré, it is possible to find a simple condition for the confinement of Dirac electrons by the moiré pattern that allows us to predict the magic angle series found numerically from a continuum model Bistritzer11 :
@@
In 2018, it was experimentally proven that this electronic localization by the moiré geometry can induce strong electronic correlations and a superconducting state for some fillings of the four flat bands Cao18a , Cao18b , DINDORKAR2023 . This new and exciting condensed matter physics joins in many aspects recent studies under the bicrystallography Gratias20 , Gratias23 , i.e., systems for which the atomic structure results from commensurable or non-commensurable modulation between two periodicities (or pseudo-periodicity), Eq. (18), of very different length (interatomic-scale periodicity and moiré-scale periodicity). In those complex systems, the consequences of crystallographic symmetries on the electronic properties are not yet well understood.
@@
Before detailing the electronic properties of flat bands in TBGs, it is interesting to note that these flat bands are not systematic when there is a moiré pattern. For example, we have not been able to obtain similar flat bands by calculations using the TB model for bilayer graphene with the similar moiré size but resulting from biaxial strain of one of the two layers (hetero-biaxial strain) without atomic relaxation. On the other hand, a very small hetero-biaxial strain (with atomic relaxation) can produce a moiré of nearly four million atoms Mesple23 . It leads to strong atomic relaxation with the AA regions almost disappearing, the appearance of swirls that delimit large AB stacking zones, and new electronic confinements in and between these zones.
@@
Obviously, as shown by the calculation of the intraband velocity $V_{TBG}({\rm K})$ at K (Fig. 6(d)), the magic angle value depends on the TB model: $\theta_{1^{st}magic}\approx 1.2^{\circ}$ and $\approx 0.9^{\circ}$ for $V_{MLG}=0.79\cdot 10^{6}$ ms-1 and $1.09\cdot 10^{6}$ ms-1 for the SK-TB model, respectively (Sec. 2.3.2). It should also be noted that $V_{\rm K}$ is very low over a fairly wide range of angles of the order of $\Delta\theta\approx 0.2^{\circ}$. This is well confirmed by the fact that the bands are fairly flat, without ever being strictly flat, for a similar range of angles around the so-called magic angle (Fig. 7). In this range of angles, other effects such as uniaxial heterostrain have a very strong effect on the bands Huder18 , Bi19 , Mesple21 (see Sec. 2.3.5).
@@
Figure 7 shows typical examples of the four flat bands for different angles around the first magic angle. Depending on the value of the angle, these bands can be isolated by minigaps or not strictly isolated, although in the latter case they are almost isolated by pseudogaps with a DOS very close to zero. So, from figure 7, for $\theta=1.248^{\circ}$, $1.203^{\circ}$ and $1.085^{\circ}$ the four flat bands are not separated by minigaps in the rest of the spectrum, whereas this is the case for $\theta=1.297^{\circ}$. It is sometimes argued in the literature that atomic relaxation is necessary to obtain these minigaps, but our results show that this is not the case, and that the shape of the bands vary strongly with angle, while remaining fairly flat overall. These bands are never perfectly flat, and it is not easy to define a “good” criterion for determining the most flat bands. Usually, as previously discussed, the intraband velocity of the Dirac cone is used, and thus the magic angle is obtained when this velocity goes to zero. However, other criteria such as bandwidth may also be relevant in some cases and may give a different value of $\theta$ for the magic angles.
@@
The local DOS in the AA zone confirms the main localization of the flat-band states in the AA zone. Detailed analysis of these flat-band states, in particular by the participation ratio calculation Trambly16 , shows that they are all equally distributed on the two layers, whereas for intermediate angles, the distribution of each state is mainly in one of the two layers (see participation ratio in Fig. 6(g)).
@@
Figure 8: Electronic properties of asymmetrically doped TBGs. TB calculation with an on-site energy difference $\Delta\epsilon$ between the two layers. (a) Sketch of the Dirac cones and total DOS of a doped rotated bilayer. (b) Energy difference $\Delta E_{D}$ between the energies of two Dirac cones in TBG versus $\Delta\epsilon$. For $\Delta\epsilon$ small, $\Delta E_{D}\approx\Delta\epsilon$ for large angles whereas $\Delta E_{D}<\Delta\epsilon$ for small angles. (c) Band velocity at the K point versus $\Delta\epsilon$. The vertical dashed lines show the value $\Delta\epsilon$ for which $\Delta E_{D}(\Delta\epsilon)$ is maximum (panel (b)). (d), (e) Bands in (25,26) TBG for $\Delta\epsilon=0$ and $\Delta\epsilon=0.2$ eV, respectively. Dashed lines are the monolayer bands. (f) Zoom of panel (e). (g) Average band velocity $V_{TBG}$ versus energy $E$ for (25,26) TBG for different $\Delta\epsilon$ values ($V_{TBG}$ is the Boltzmann velocity computed from the velocity operator and Eq. (76)). TBG: $(5,9)$ $\theta=18.73^{\circ}$, $(2,3)$ $\theta=13.17^{\circ}$, $(3,4)$ $\theta=9.43^{\circ}$, $(6,7)$ $\theta=5.08^{\circ}$, $(8,9)$ $\theta=3.89^{\circ}$, $(12,13)$ $\theta=2.13^{\circ}$, $(25,26)$ $\theta=1.297^{\circ}$. SK-TB model is the one which results in an intraband monolayer velocity $V_{MLG}=0.79\cdot 10^{6}$ ms-1 (Sec. 2.3.2). TBGs are rigid, i.e., without atomic relaxation. From Trambly16 .
We now consider asymmetrically doped TBG, i.e., TBG with an on-site energy difference $\Delta\epsilon=\epsilon({\rm layer\,1})-\epsilon({\rm layer\,2})$ between the two layers (Fig. 8). It may be due to asymmetrical doping of the two graphene monolayers (e.g., from substrate doping) or to the action of an external electric field perpendicular to the plane, which can be controlled by a gate. For large twist angles $\theta$, when the states of layer 1 and layer 2 are decoupled, the asymmetric doping simply shifts the two Dirac cones by $\Delta\epsilon$, so the difference in the two Dirac energies is $\Delta E_{D}\approx\Delta\epsilon$ for $\Delta\epsilon$ small. For smaller $\theta$, intermediate angles $\theta$ with $\theta\gtrsim 3^{\circ}$, the main effect of a small $\Delta\epsilon$ is to bring the energies of the Dirac cones of each layer closer together (so $\Delta E<\Delta\epsilon$) thus increasing the reduction in electron velocity Cherkez15 , Trambly16 (Fig. 8(b,c)). For smaller angles close to the magic angle, $\Delta\epsilon$ shifts the band energies and modifies them, but without making flatness disappear and the sharp reduction in band velocity is still very much present (Fig. 8(e,f,g)).
@@ 2.3.4 Quantum diffusion
Figure 9: Diffusivity $\mathcal{D}$ versus scattering time $\tau$ in graphene, TBGs (5,9), (6,7), and (25,26), with rotation angle $\theta=18.73^{\circ}$, $5.08^{\circ}$ and $1.297^{\circ}$, respectively. For (25,26) TBG the Boltzmann (B) term and non-Boltzmann (NB) term are shown, $\mathcal{D}(E_{F},\tau)=\mathcal{D}_{\rm B}(E_{F},\tau)+\mathcal{D}_{\rm NB}(E_{F},\tau)$ (see Eq. (79)). These calculations are performed with an energy broadening of 10 meV around $E_{F}$. For (25,26) TBG that corresponds mainly to the flat bands (Fig. 7(a)), i.e., the minimum of $V_{\mathrm{B}}$ (Fig. 8(g)). (a) $\Delta\epsilon=0$, $E_{F}=0.12$ eV (see bands and DOS in Fig.7(a)), and (b) $\Delta\epsilon=0.2$ eV, $E_{F}=-0.1$ eV. $\Delta\epsilon$ the energy difference between the on-site energy in the two layers (Sec. 2.3.3). SK-TB model is the one which results in an intraband monolayer velocity $V_{MLG}=0.79\cdot 10^{6}$ ms-1 (Sec. 2.3.2). TBGs are rigid, i.e., without atomic relaxation. For the numerical method see A.2.1. From Trambly16 .
@@
In this section, we describe the consequences of the particular localization of TBG flat band states on their quantum diffusion properties Trambly16 , Andelkovic18 , Sinha24 , guerrero24_tBLG_disorder_Kubo , GobboKhun25 . Here we are not considering the defect effect, which produces particular states, and will be studied in the next section (Sec. 3.2.3). We are only considering quantum diffusion in the defect-free medium, and we are treating the effect of disorder by a scattering time $\tau$. As explained in section 2.1 (see also A.2), the diffusivity $\mathcal{D}$ at every energy $E$ and scattering time $\tau$ is the sum of a Boltzmann term, $\mathcal{D}_{\rm B}(E_{F},\tau)=V_{\mathrm{B}}^{2}\tau$, and a non-Boltzmann term $\mathcal{D}_{\rm NB}$. For large $\tau$, $\mathcal{D}_{\rm NB}$ decreases when $\tau$ increases, and $\mathcal{D}_{\rm NB}\rightarrow 0$ when $\tau\rightarrow+\infty$. Thus in crystals, $\mathcal{D}\approx\mathcal{D}_{\rm B}$ when $\tau\rightarrow+\infty$.
@@
The diffusivity, calculated by diagonalization (A.2), for monalayer graphene and several cases of TBGs are presented in Fig. 9 for different $E_{F}$ values and for doped or undoped bilayers. For graphene and bilayers with large and intermediate rotation angles, $\mathcal{D}\approx\mathcal{D}_{\rm B}$ at every energy. The only effect of the non-Boltzmann term is a change in the slope of $\mathcal{D}(\tau)$ at scattering time $\tau\approx\hbar/E$ as explained in the appendix of Ref. Trambly16 . Eventually at small scattering time, $\tau\ll\hbar/E$, the interband transition between the two bands of each Dirac cone contributes significantly. In the case of graphene, with a first-neighbor coupling Hamiltonian, the non-Boltzmann term $\mathcal{D}_{\rm NB}$ is equal to the Boltzmann term $\mathcal{D}_{\rm B}$ and then $\mathcal{D}=2\mathcal{D}_{\rm B}$ for $\tau\ll\hbar/E$ Trambly16 . This effect is related to the phenomenon of jittery motion, also called Zitterbewegung Castro09_RevModPhys , which is important in the optical conductivity. In graphene and bilayers with large rotation angles, it occurs for very small scattering time values that are too small to be significant experimentally. However, in rotated bilayers with very small rotation angle $\theta$ close to the first magic angle, for states at energy where velocity is very small (i.e., energies close to the Dirac energy), the Boltzmann terms in the transport coefficients (see equations (8) and in A.2) decrease and the non-Boltzmann terms become significant. As we have seen in the case of quasiperiodic or approximant structures, this has a strong effect on the charge carrier diffusion regime and therefore on the effect of disorder.
@@
For instance, Fig. 9 shows that for (25,26) bilayers ($\theta=1.297^{\rm o}$), the diffusivity $\mathcal{D}$ is strongly affected by the non-Boltzmann term for realistic $\tau$ values Wu07 . It results in a smaller diffusivity with respect to the case of graphene, that is not at all ballistic; indeed $\mathcal{D}$ is not at all linear as a function of $\tau$ for realistic values of the scattering time, $\tau\approx 10^{-14}$$10^{-13}$s. Even more, at charge neutrality (i.e., $E_{F}=0.012$ eV for (25,26) TBG see Fig. 7(a)) and without asymmetric doping between the two layers (Fig. 9(a)), diffusivity decreases from $\mathcal{D}=6\cdot 10^{-3}$ m2s-1 to $3\cdot 10^{-3}$ m2s-1, as $\tau$ increases from $\tau\approx 4\cdot 10^{-15}$ s and $10^{-13}$ s. This decrease in conductivity by a factor of 2 due to increased disorder, which may be static (defects) or dynamic (temperature) clearly shows that disorder can lead to an increase in conductivity of the flat-band states, as expected in the sub-diffusive regime (see Sec. 2.2.2). Since flat-band states are localized over large areas of space (several hundreds atoms in AA zones) and their band velocity is very low, it is clear that these states satisfy the SVR condition (Eq. (6)). It also shows indirectly that flat-band states are fragile and can be destroyed by disorder. Recent real space calculations of conductivity using Kubo’s formula have shown such behavior for a magic-angle TBG in the presence of Anderson disorder guerrero24_tBLG_disorder_Kubo . In asymmetric doped TBG (Fig. 9(b)), similar effects occur at energies with small Boltzmann velocity (Fig. 8(g)). This regime, called small velocity regime (SVR), where non-Boltzmann terms dominate transport properties, has already been observed in systems with complex atomic structure such as quasicrystals and complex metallic alloys (Sec. 2.2.1). Let us recall that SVR is reached when the mean free path $V\tau$ of charge carriers is smaller than the spatial extent $L_{w}$ of the corresponding wave packet for large time $t$, i.e., $t\gtrsim\tau$.